One of your customers is delinquent on his accounts payable balance. You’ve mutually agreed to a repayment schedule of $750 per month. You will charge 1.8 percent per month interest on the overdue balance.

If the current balance is $17,000, how long will it take for the account to be paid of
Answer #1

Here we need to find the length of an annuity. We know the interest rate, the PV, and the payments. Using the PVA equation:

PVA = C({1 – [1/(1 + r)t]} / r)

$17,000 = $750{[1 – (1/1.0180)t] / 0.0180}

Now we solve for t:

1/1.0180t = 1 − {[($17,000) / ($750)](0.0180)}

1/1.0180t = 0.5920

1.0180t = 1 / (0.5920) = 1.689

t = ln 1.689 / ln 1.0180 = 29.39 months